Sol Summary – February 12th

Crew 205 Sol Summary Report 12- FEB – 2019

Sol: 2

Summary Title: Adapting to Mars

Author’s name: Natalia Larrea

Mission Status: Nominal

Sol Activity Summary: We started the day with what is becoming our daily routine: waking up at 7 am, having coffee and breakfast together and by 8.30 am start our work. In the morning, the first EVA team set off on rovers Spirit and Opportunity Northwards on Cow Dung Road towards the “Moon” areas. They passed Galileo road (losing contact with MDRS as expected shortly after). Before reaching Cowboy Corner, Spirit reported having 75% SOC, Spirit’s battery reduced rapidly and the rovers were stopped between Tank Wash and White Moon with Spirit on 63% and Opportunity at 75%. The EVA was modified to be completed on foot by only visiting, exploring and sampling areas within White Moon and the easternmost fringes of Beige Moon. The turnings of the road that enter to the Moon were recorded on GPS for future use in EVAs. Some excellent samples of quartz and sulphur containing deposits were collected and will be analysed in the coming days before being described fully in the upcoming Science Report. Meanwhile, in the Hab, the rest of the team continue working in the GreenHab daily duties, EVA planning and astronomy projects. Lunch was ready for the returning EVA team and the crew ate together in good spirits.

Equipment was prepared for use in the second EVA of the day, including a drone with a vertical camera and the ability to connect to a laptop for images to be reviewed in the field. As planned, the EVA team started getting prepared for its mission to North Ridge. However, while preparing, it was found that some of the suits were not fully functional and others, were still charging from the EVA in the morning. The team initially thought only of the suits was to be fully charged and operational. For this reason, putting the safety of the crew as the top priority, the commander took the decision to cancel the EVA. After the fact, it was found that a second spacesuit was also fully charged. All crew members thus, stayed in the Hab working “from home” in their different duties and projects. Tonight, following our daily crew meeting, we plan to enjoy a “Martian-Spanish” tapas dinner.

Look Ahead Plan: Tomorrow we plan to continue with the initial plan of today. A first EVA team will head in the morning will head to Maxwell Montes. A second team will head the Old Repeater Point and North Ridge to test the drone. GreenHab, astronomy and outreach activities will continue as usual.

Anomalies in work: EVA#04 (in the afternoon) was cancelled (please refer to EVA report and description above for further information)

Weather: Clear skies all day long

Crew Physical Status: Nominal


EVA#03: 4254500/516500 (Gray Moon, not reached), 4254350/516500 (Beige Moon), 4254550/517550 (White Moon)

EVA#04: No conducted (please refer to the EVA report and description above for more information)

Reports to be filed: Sol Summary, EVA request, Operations Report, Astronomy Report, Greenhab Report, Journalist Report

Support Requested: Some spacesuits not working properly and we are missing a charger for one of them (please refer to operations report for further information).

Copyright © The Mars Society. All rights reserved. | Main Site